find radius of curvature of
16y^2=4x^4-x^6

Answer by Alan3354(69443)

find radius of curvature of
16y^2=4x^4-x^6
--------------------
Find dy/dx and d^2y/d^2x (y' and y")
---
16y^2=4x^4-x^6
32ydy = 16x^3dx - 6x^5dx
y' = (8x^3 - 3x^5)/16y
-------
16ydy = 8x^3dx - 3x^5dx
16d^2y = 24x^2d^2x - 15x^4d^2x
y" = (24x^2 - 15x^4)/16
===========================
Curv = (1 + y"^2)^(3/2)/y"
-----------
Pick a point and sub the values.